### v2

# Approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search

Abstract
We propose a variant of the approximate Bregman proximal gradient (ABPG) algorithm for minimizing the sum of a smooth nonconvex function and a nonsmooth convex function. ABPG is known to converge globally to a stationary point even when the smooth part of the objective function does not have a globally Lipschitz continuous gradient, and its iterates can often be expressed in closed form. However, ABPG relies on an Armijo line search to guarantee global convergence, which can slow down its practical performance. To address this issue, we propose a variant of ABPG with a variable metric Armijo–Wolfe line search. Under the variable metric Armijo–Wolfe condition, we establish global subsequential convergence of the algorithm. Moreover, assuming the Kurdyka–Łojasiewicz property, we also prove that the algorithm globally converges to a stationary point. Numerical experiments on $\ell_{p}$-regularized least squares problems and nonnegative linear inverse problems demonstrate that the proposed algorithm outperforms existing algorithms.
keywords:
Composite nonconvex nonsmooth optimization, Bregman proximal gradient algorithms, Kurdyka–Łojasiewicz property, Bregman divergence
1 Introduction
We consider composite nonconvex optimization problems of the form
$\min_{x\in\operatorname{cl}C}\quad\Psi(x)\coloneqq f(x)+g(x),$ (1.1)
where $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a continuously differentiable function, $g:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a possibly nondifferentiable convex function, and $\operatorname{cl}C$ is the closure of a nonempty open convex set $C\subset\mathbb{R}^{n}$. Optimization problems of the form (1.1) arise in various applications, including the maximum a posteriori (MAP) estimate [MAP, MAP2], ridge regression [ridge], the least absolute shrinkage and selection operator (LASSO) [LASSO]. In machine learning and signal processing, regularization or penalty terms are often introduced to prevent overfitting and impose the model structure. Some regularization terms are not necessarily differentiable.
Numerous algorithms using proximal mappings have been proposed to solve (1.1). For instance, the proximal gradient method [Bruck1975-et, lions1979splitting, PASSTY1979383] and the fast iterative shrinkage-thresholding algorithm (FISTA) [fista] belong to the class of proximal algorithms. Convergence analysis of these algorithms with constant step-sizes typically relies on the global Lipschitz continuity of $\nabla f$, i.e., there exists $L>0$ such that $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for any $x,y\in\mathbb{R}^{n}$. This condition is often restrictive and does not hold in certain applications in signal processing and machine learning.
Bolte et al. [Bolte-2018] proposed the Bregman proximal gradient algorithm (BPG). This algorithm globally converges under the smooth adaptable property [Bolte-2018], also called relative smoothness [Lu-haihao], which is a relaxation of the global Lipschitz continuity of $\nabla f$. In recent years, the Bregman proximal gradient method has been improved from various perspectives. Hanzely et al. [Hanzely2021-sz] proposed accelerated Bregman proximal gradient algorithms for convex optimization problems using the triangle scaling property. Mukkamala et al. [Mukkamala2020-ea] proposed an accelerated version of BPG. Some researchers have applied Bregman proximal-type algorithms to linear inverse problems [Bauschke2017-hg, Takahashi], nonnegative matrix factorization [Mukkamala2019-mk, Takahashi2026-rv], and blind deconvolution [Takahashi2023-uh].
Since the subproblem of BPG cannot always be solved in closed form and is sometimes hard to solve depending on the Bregman distance, Takahashi and Takeda [Takahashi] proposed the approximate Bregman proximal gradient algorithm (ABPG), whose subproblem is easier to solve. Instead of the Bregman distance, ABPG uses the approximate Bregman distance (see also (3.2)), which is the second-order approximation of the Bregman distance. The subproblem of ABPG can be written by the sum of a quadratic function and a regularizer. Moreover, if the Bregman distance is separable, the subproblem of ABPG is reduced to $n$ independent one-dimensional optimization problems. ABPG uses the line search procedure to ensure the accuracy of the approximate Bregman distance. However, the global convergence of ABPG has not been established when $g\not\equiv 0$, and line search procedures can lead to slow convergence in practice.
In this paper, we propose a new algorithm, named the approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search (ABPG-VMAW). The line search procedure of this algorithm is inspired by variable metric inexact line search based algorithms [Bonettini2016-bn, Bonettini_2017]. In the same way as ABPG, the subproblem of ABPG-VMAW is defined by
$y^{k}=\operatorname*{argmin}_{u\in\operatorname{cl}C}\left\{\langle\nabla f(x^{k}),u-x^{k}\rangle+g(u)+\frac{1}{\lambda}\tilde{D}_{\phi}(u,x^{k})\right\},$ (1.2)
where $x^{k}\in\operatorname{cl}C$, $\lambda>0$, and $\tilde{D}_{\phi}(u,x):=\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle$ is the approximate Bregman distance with a twice continuously differentiable convex function $\phi$. The search direction of ABPG is defined by $d^{k}=y^{k}-x^{k}$, and ABPG searches for $t_{k}\in(0,1]$ in each iteration to decide the step-size. The Armijo-like condition adopted in [Takahashi] is so stringent that computing $t_{k}$ can be time-consuming, and it may force $t_{k}$ to be very small, which causes slow convergence. In this paper, we adopt a relaxed condition that allows larger values of $t_{k}$. In addition to this, inspired by the Armijo–Wolfe-like condition, which Lewis and Overton [Lewis2013] applied to the quasi-Newton methods, we also propose a curvature condition for proximal algorithms. It aims to avoid excessively small step-sizes while ensuring that the search direction $d^{k}$ approaches 0 as $k\to\infty$. Similar to Bonettini et al. [Bonettini_2017], we also add a rule at the end of each iteration to select, as the updated point, the one that yields a smaller value of the objective function between $y^{k}$ and the point provided from the line search.
Through these modifications, we establish that, under standard assumptions, accumulation points of a sequence generated by ABPG-VMAW are stationary points. Furthermore, by assuming the Kurdyka–Łojasiewicz property [90074980-9c0c-33c4-b119-70171dab0b45] for $\Psi$, we prove that our algorithm achieves global convergence even for $g\not\equiv 0$.
Moreover, numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems demonstrate that ABPG-VMAW outperforms ABPG and other existing algorithms. In particular, the reduction of the objective function value within a small number of iterations is faster for ABPG-VMAW than for ABPG.
The structure of this paper is as follows. Section 2 introduces essential notation such as the subdifferential, the Bregman distances, and the Kurdyka–Łojasiewicz property. In Section 3, we propose ABPG-VMAW and discuss its line search conditions. Section 4 presents properties of ABPG-VMAW and its global convergence. Section 5 presents numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems. Finally, in Section 6, we present conclusions and future research directions.
Notation
In this paper, we use the following notation. Let $\mathbb{R}$ and $\mathbb{R}_{+}$ be the set of real numbers and nonnegative real numbers, respectively. Let $\mathbb{R}^{n}$ and $\mathbb{R}_{+}^{n}$ be the real space of $n$ dimensions and the nonnegative orthant of $\mathbb{R}^{n}$, respectively. Let $\mathbb{R}^{n\times m}$ be the set of $n\times m$ real matrices. The identity matrix is $I\in\mathbb{R}^{n\times n}$. Let $|x|$ and $x^{p}$ be the elementwise absolute and $p$th power vectors of $x\in\mathbb{R}^{n}$, respectively. Given a real number $p\geq 1$, the $\ell_{p}$ norm is defined by $\|x\|_{p}=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{1/p}$. Let $\lambda_{\max}(M)$ be the largest eigenvalue of a symmetric matrix $M\in\mathbb{R}^{n\times n}$.
Let $B(\bar{x},r)=\{x\in\mathbb{R}^{n}\,|\,\|x-\bar{x}\|\leq r\}$ denote the ball with center $\bar{x}\in\mathbb{R}^{n}$ and radius $r>0$. Let $\operatorname{int}C$ and $\operatorname{cl}C$ be the interior and the closure of a set $C\subset\mathbb{R}^{n}$, respectively. The distance from a point $x\in\mathbb{R}^{n}$ to $C$ is defined by $\operatorname{dist}(x,C):=\inf_{y\in C}\|x-y\|$. The indicator function $\delta_{C}$ is defined by $\delta_{C}(x)=0$ for $x\in C$ and $\delta_{C}(x)=+\infty$ otherwise. The sign function $\operatorname{sgn}(x)$ is defined by $\operatorname{sgn}(x)=-1$ for $x<0$, $\operatorname{sgn}(x)=0$ for $x=0$, and $\operatorname{sgn}(x)=1$ for $x>0$.
Given $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}$, we define a set $\Psi(y)<\Psi<\Psi(y)+z]$ as the set of all $x$ in the subset of $\mathbb{R}^{n}$ that satisfy $\Psi(y)<\Psi(x)<\Psi(y)+z$. Given $k\in\mathbb{N}$, let $\mathcal{C}^{k}$ be the class of $k$-times continuously differentiable functions.
2 Preliminaries
2.1 Subdifferentials
First, we introduce the definitions of subdifferentials. For an extended-real-valued function $f:\mathbb{R}^{n}\rightarrow[-\infty,+\infty]$, the effective domain of $f$ is defined by $\operatorname{dom}f\coloneqq\{x\in\mathbb{R}^{n}\,|\,f(x)<+\infty\}$. The function $f$ is proper if $f(x)>$-$\infty$ for all $x\in\mathbb{R}^{n}$ and $\operatorname{dom}f\neq\emptyset$.
Definition 2.1 (Regular and Limiting Subdifferentials [rockafellar2009variational, Definition 8.3]).
Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function.
  1. (i)
    The regular subdifferential of $f$ at $x\in\operatorname{dom}f$ is defined by
    $\hat{\partial}f(x)=\left\{\xi\in\mathbb{R}^{n}\,\middle|\,\liminf_{y\to x,y\neq x}\frac{f(y)-f(x)-\langle\xi,y-x\rangle}{\|x-y\|}\geq 0\right\}.$ (2.1)
    When $x\not\in\operatorname{dom}f$, we set $\hat{\partial}f(x)=\emptyset$.
  2. (ii)
    The limiting subdifferential of $f$ at $x\in\operatorname{dom}f$ is defined by
    $\partial f(x)=\left\{\xi\in\mathbb{R}^{n}\,\middle|\,\exists x^{k}\xrightarrow{f}x,\xi^{k}\to\xi,\forall k\in\mathbb{N},\xi^{k}\in\hat{\partial}f(x^{k})\right\},$ (2.2)
    where $x^{k}\xrightarrow{f}x$ means $x^{k}\to x$ and $f(x^{k})\to f(x)$.
Generally, $\hat{\partial}f(x)\subset\partial f(x)$ holds for all $x\in\mathbb{R}^{n}$ [rockafellar2009variational, Theorem 8.6]. We define $\operatorname{dom}\partial f\coloneqq\{x\in\mathbb{R}^{n}|\partial f(x)\neq\emptyset\}$. If $f$ is convex, the regular and limiting subdifferentials coincide with the (classical) subdifferential [rockafellar2009variational, Proposition 8.12].
For a proper and convex function $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$, the directional derivative of $f$ at $x\in\operatorname{dom}f$ in the direction $d$ is given by
$f^{\prime}(x;d)=\lim_{t\to+0}\frac{f(x+td)-f(x)}{t}.$ (2.3)
From [Rockafellar1970-if, Theorem 23.1], $\frac{f(x+td)-f(x)}{t}$ is monotonically non-decreasing with respect to $t$ for $t>0$. The limit on the right-hand side always exists if $\pm\infty$ is allowed as a possible limit value. For any $x\in\operatorname{dom}f$, $\xi\in\partial f(x)$ if and only if $f^{\prime}(x;d)\geq\langle\xi,d\rangle$ holds for any $d\in\mathbb{R}^{n}$ [Rockafellar1970-if, Theorem 23.2].
2.2 Bregman Distances
Let $C$ be a nonempty and convex subset of $\mathbb{R}^{n}$. We introduce the kernel generating distance [Bolte-2018] and the Bregman distance.
Definition 2.2 (Kernel Generating Distances [Bolte-2018, Definition 2.1]).
A function $\phi:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is called a kernel generating distance associated with $C$ if it satisfies the following conditions:
  1. (i)
    $\phi$ is a proper, lower semicontinuous, and convex function, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$.
  2. (ii)
    $\phi$ is $\mathcal{C}^{1}$ on $\operatorname{int}\operatorname{dom}\phi\equiv C$.
We denote $\mathcal{G}(C)$ as the class of kernel generating distances associated with $C$.
Definition 2.3 (Bregman Distances [Bregman1967TheRM]).
For a kernel generating distance $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\rightarrow\mathbb{R}_{+}$ is defined by
$D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle.$ (2.4)
Because the Bregman distance does not satisfy the symmetry and the triangle inequality, it is not a distance. Due to the convexity of $\phi$, $D_{\phi}(x,y)\geq 0$ for any $(x,y)\in\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi$. If $\phi$ is strictly convex, $D_{\phi}(x,y)=0$ holds if and only if $x=y$. We also show some examples of Bregman distances.
Example 2.4.
  • Mahalanobis distance: Let $\phi(x)=\frac{1}{2}\langle Ax,x\rangle$ for a positive definite matrix $A\in\mathbb{R}^{n\times n}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$. Then, we have $D_{\phi}(x,y)=\frac{1}{2}\langle A(x-y),x-y\rangle$, which is called the Mahalanobis distance. When $A=I$, the Mahalanobis distance corresponds with the squared Euclidean distance, i.e., $D_{\phi}(x,y)=\frac{1}{2}\|x-y\|^{2}$.
  • Kullback–Leibler divergence [Kullback1951-yv]: Let $\phi$ be the Boltzmann–Shannon entropy, i.e., $\phi(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ with $0\log 0=0$ and $\operatorname{dom}\phi=\mathbb{R}_{+}^{n}$. Then, we have $D_{\phi}(x,y)=\sum_{i=1}^{n}x_{i}\log\frac{x_{i}}{y_{i}}$, which is called the Kullback–Leibler divergence.
  • Itakura–Saito divergence [Itakura1968-en]: Let $\phi$ be the Burg entropy, i.e., $\phi(x)=-\sum_{i=1}^{n}\log x_{i}$ and $\operatorname{dom}\phi=\mathbb{R}^{n}_{++}$. Then, we have $D_{\phi}(x,y)=\sum_{i=1}^{n}\left(\frac{x_{i}}{y_{i}}-\log\frac{x_{i}}{y_{i}}-1\right)$, which is called the Itakura–Saito divergence.
See [Bauschke2017-hg, Bauschke1997-vk, Lu-haihao] and [Dhillon2008-zz, Table 2.1] for more examples.
2.3 Kurdyka–Łojasiewicz Property
The Kurdyka–Łojasiewicz (KL) property is an essential assumption to establish global convergence. Attouch et al. [90074980-9c0c-33c4-b119-70171dab0b45] extended the Łojasiewicz gradient inequality [Kurdyka1998, AIF_1993__43_5_1575_0] to nonsmooth functions.
For $v>0$, we define $\Xi_{v}$ as a set of all continuous concave functions $\psi:[0,v)\rightarrow\mathbb{R}_{+}$ that are $\mathcal{C}^{1}$ on $(0,v)$ and satisfies $\psi(0)=0$, and whose derivative $\psi^{\prime}(x)$ is positive on $(0,v)$. We define the Kurdyka–Łojasiewicz property.
Definition 2.5 (Kurdyka–Łojasiewicz Property [90074980-9c0c-33c4-b119-70171dab0b45]).
Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function. The function $f$ is said to satisfy the Kurdyka–Łojasiewicz property (for short: KL property) at $\bar{x}\in\operatorname{dom}\partial f$ if there exist $v\in(0,+\infty]$, a neighborhood $U$ of $\bar{x}$, and a function $\psi\in\Xi_{v}$, such that for any $x\in U\cap[f(\bar{x})<f<f(\bar{x})+v]$, the following inequality holds:
$\psi^{\prime}(f(x)-f(\bar{x}))\operatorname{dist}(0,\partial f(x))\geq 1.$ (2.5)
Moreover, $f$ is called a KL function if $f$ satisfies the KL property at each point of $\operatorname{dom}\partial f$.
The uniformized KL property is established by the KL property.
Lemma 2.6 (Uniformized KL property [bolte-2014, Lemma 6]).
Assume that $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is a proper and lower semicontinuous function. If $f$ takes a constant value on some compact set $\Gamma$, and satisfies the KL property on $\Gamma$, then there exist $v,\epsilon\in(0,+\infty]$, and $\psi\in\Xi_{v}$ such that, for any $\bar{x}\in\Gamma$, and any $x\in\mathbb{R}^{n}$ satisfying $\operatorname{dist}(x,\bar{x})<\epsilon$ and $x\in[f(\bar{x})<f<f(\bar{x})+v]$, the following inequality holds:
$\psi^{\prime}(f(x)-f(\bar{x}))\operatorname{dist}(0,\partial f(z))\geq 1.$ (2.6)
3 Proposed Algorithm: Approximate Bregman Proximal Gradient Algorithm with Variable Metric Armijo–Wolfe Line Search
Throughout this paper, we make the following assumptions.
Assumption 3.1.
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}C=\operatorname{cl}\operatorname{dom}\phi$ is $\mathcal{C}^{2}$ on $C=\operatorname{int}\operatorname{dom}\phi$.
  2. (ii)
    $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$ and $\mathcal{C}^{1}$ on $C$.
  3. (iii)
    $g:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ is proper, lower semicontinuous, and convex with $C\subset\operatorname{dom}g$.
  4. (iv)
    $\Psi^{*}\coloneqq\inf_{x\in\operatorname{cl}C}\Psi(x)>$-$\infty$.
  5. (v)
    For any $x\in\operatorname{int}\operatorname{dom}\phi$ and $\lambda>0$, $u\mapsto\lambda g(u)+\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle$ is supercoercive, that is,
    $\lim_{\|u\|\to\infty}\frac{\lambda g(u)+\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle}{\|u\|}=\infty.$ (3.1)
Assumption 3.1(i)-(iv) are standard assumptions for Bregman-type algorithms [Bolte-2018, Takahashi] and are generally satisfied in practice. For any $x\in C$, $\partial(g+\delta_{\operatorname{cl}{C}})(x)=\partial g(x)+\partial\delta_{\operatorname{cl}{C}}(x)=\partial g(x)$ holds because $x$ is an interior point of $C$ and $\operatorname{dom}g$ from Assumption 3.1(iii) and $\partial\delta_{\operatorname{cl}{C}}(x)=\{0\}$. For example, Assumption 3.1(v) holds if $\phi$ is strongly convex. Note that we will assume the strong convexity of $\phi$ in Assumption 4.1.
3.1 Approximate Bregman Proximal Gradient Algorithm
Let $\phi\in\mathcal{G}(C)$ be $\mathcal{C}^{2}$ on $C$. Takahashi and Takeda [Takahashi] define the approximate Bregman distance $\tilde{D}_{\phi}(u,x)\geq 0$, using a second-order approximation of $\phi(u)$ for $u\in\operatorname{dom}\phi$ around point $x\in\operatorname{int}\operatorname{dom}\phi$, as
$\tilde{D}_{\phi}(u,x)\coloneqq\frac{1}{2}\langle\nabla^{2}\phi(x)(u-x),u-x\rangle\simeq D_{\phi}(u,x).$ (3.2)
Note that $D_{\phi}(u,x)\leq\tilde{D}_{\phi}(u,x)$ or $D_{\phi}(u,x)\geq\tilde{D}_{\phi}(u,x)$ does not necessarily hold for any $x$ and $u$. Therefore, a line search was incorporated into the proposed algorithm.
The Bregman proximal gradient mapping [Bolte-2018] at a point $x\in C$ for a parameter $\lambda>0$ is defined by
$\mathcal{T}_{\lambda}(x)\coloneqq\operatorname*{argmin}_{u\in\operatorname{cl}C}\left\{\langle\nabla f(x),u-x\rangle+g(u)+\frac{1}{\lambda}D_{\phi}(u,x)\right\}.$ (3.3)
Instead of (3.3), the approximate Bregman proximal gradient mapping [Takahashi] at a point $x\in C$ is defined by
$\tilde{\mathcal{T}}_{\lambda}(x)\coloneqq\operatorname*{argmin}_{u\in\operatorname{cl}C}\left\{\langle\nabla f(x),u-x\rangle+g(u)+\frac{1}{\lambda}\tilde{D}_{\phi}(u,x)\right\}.$ (3.4)
Using Assumption 3.1(iii) and the positive semidefiniteness of $\nabla^{2}\phi$, (3.4) is a convex optimization problem.
Assumption 3.2.
For any $x\in C$ and any $\lambda>0$, $\tilde{\mathcal{T}}_{\lambda}(x)\subset C$ holds.
Assumption 3.2 ensures that the points generated by ABPG-VMAW are feasible. Obviously, when $C\equiv\mathbb{R}^{n}$, Assumption 3.2 holds. In the same discussion as [Bolte-2018, p. 2136] and [Takahashi, p.235], if $\phi$ is strongly convex, the envelope function $\inf_{u\in\operatorname{cl}{C}}\left\{\langle\nabla f(x),u-x\rangle+g(u)+\frac{1}{\lambda}\tilde{D}_{\phi}(u,x)\right\}$ is prox-bounded from [rockafellar2009variational, Exercise 1.24]. We have the following well-posedness result.
Lemma 3.3 (Well-posedness of $\tilde{\mathcal{T}}_{\lambda}$ [Takahashi, Lemma 12]).
Suppose that Assumptions 3.1 and 3.2 hold. For any $x\in\operatorname{int}\operatorname{dom}\phi$ and any $\lambda>0$, the approximate Bregman proximal gradient mapping $\tilde{\mathcal{T}}_{\lambda}(x)$ is a nonempty compact subset of $C$.
3.2 Variable Metric Armijo–Wolfe Line Search
Since $\Psi(y^{k})\leq\Psi(x^{k})$ is not necessarily guaranteed for the solution of the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$ with any $\lambda>0$, Takahashi and Takeda [Takahashi] introduced the line search procedure for ABPG. To ensure global convergence, we improved the condition of the line search procedure. The new condition is inspired by the line search method based on the Armijo–Wolfe condition proposed by Lewis and Overton [Lewis2013] and Miantao et al. [Miantao2024], and on the Armijo-like line search method introduced by Bonettini et al. [Bonettini_2017]. We also execute an update step to take a point corresponding to a smaller value of the objective function at the end of each iteration, inspired by Bonettini et al. [Bonettini_2017].
In order to define the search direction $d^{k}=y^{k}-x^{k}$, we solve the subproblem $y^{k}\in\tilde{\mathcal{T}}_{\lambda}(x^{k})$. Let $0<c_{1}<c_{2}<1$ and $\xi^{k}\in\partial g(x^{k})$. To ensure that $\Psi(x^{k+1})$ is sufficiently smaller than $\Psi(x^{k})$, we impose the following condition on $t$:
$\Psi(x^{k}+td^{k})+\delta_{\operatorname{cl}{C}}(x^{k}+td^{k})$ (3.5)
$\qquad<\Psi(x^{k})+c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right).$ (3.6)
To ensure that $x^{k+1}\in C$, we use $\delta_{\operatorname{cl}{C}}(x^{k}+td^{k})$. Furthermore, to avoid excessively small step-sizes $t_{k}>0$, which could slow down convergence, we impose the condition given by
$\langle\nabla f(x^{k}+td^{k})+\xi^{k},d^{k}\rangle>c_{2}\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle.$ (3.7)
We consider $t_{k}$ as $t$ satisfying both (3.6) and (3.7) simultaneously. Here, by rearranging the inequalities of the line search procedure, we define
$A_{k}(t)$ $\coloneqq\Psi(x^{k}+td^{k})+\delta_{\operatorname{cl}{C}}(x^{k}+td^{k})-\Psi(x^{k})$ (3.8)
$\quad-c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right),$ (3.9)
$W_{k}(t)$ $\coloneqq\langle\nabla f(x^{k}+td^{k})+\xi^{k},d^{k}\rangle-c_{2}\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle.$ (3.10)
The line search conditions (3.6) and (3.7) can be rewritten as $A_{k}(t)<0$ and $W_{k}(t)>0$, respectively.
Now we are ready to describe the proposed algorithm and its line search procedure for solving (1.1). The subproblem $\tilde{\mathcal{T}}_{\lambda}(x^{k})$ on line 1 is convex, and it is strongly convex if $\phi$ is strongly convex (see also Assumption 4.1). To obtain the step-size $t_{k}$ satisfying $A_{k}(t_{k})<0$ and $W_{k}(t_{k})>0$ on line 1, we can use, for example, the bisection method (see, for more details, Section A). Moreover, $A_{k}(t_{k})<0$ implies $x^{k}+t_{k}d^{k}\in\operatorname{cl}{C}$ because of the term $\delta_{\operatorname{cl}{C}}(x^{k}+t_{k}d^{k})$ of $A_{k}(t_{k})$.
Remark 3.4 (Comparison with existing line search conditions).
For simplicity, we assume $C=\mathbb{R}^{n}$, so that $\delta_{\operatorname{cl}{C}}\equiv 0$. Our variable metric line search condition (3.6) allows a larger $t$ than the classical Armijo condition:
$\Psi(x^{k}+td^{k})$ (3.11)
$\qquad<\Psi(x^{k})+c_{1}t(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k}))$ (3.12)
$\qquad\leq\Psi(x^{k})+c_{1}t\left(\langle\nabla f(x^{k}),d^{k}\rangle+g(x^{k}+d^{k})-g(x^{k})+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\right),$ (3.13)
where the second inequality follows from the variable metric $\frac{1}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\geq 0$. When $\phi$ is strongly convex, the above second inequality holds with strict inequality for $d^{k}\neq 0$. Moreover, the parameter $\lambda$ can be chosen freely. As a practical guideline for choosing $\lambda$, we refer the reader to Remark 3.5.
On the other hand, our curvature condition (3.7), which uses $\xi^{k}$ on both sides, differs from the existing one. Although this may seem somewhat odd at first glance, it will play an important role in the proof of Lemma 4.9. When $g$ is nonsmooth, its subgradient typically fails to satisfy standard regularity conditions such as Lipschitz continuity, which is why $\xi^{k}$ appears on both sides.
Although we can choose any $\lambda>0$, it is better to use $\lambda<1/L$ for some $L$ as follows (see, for specific examples, Section 5).
Remark 3.5.
As already mentioned in Remark 3.4, the parameter $\lambda$ can be chosen to be any positive scalar. We now provide a practical guideline for selecting $\lambda$ to reduce the number of iterations. In practice, it is better to choose $\lambda<1/L$, where $L>0$ is a parameter given by the smooth adaptable property, i.e., the pair $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad[Bolte-2018] if there exists $L>0$ such that both $L\phi-f$ and $L\phi+f$ are convex on $C$. The $L$-smad property provides the first-order approximation of $f$ by its descent lemma [Bolte-2018, Lemma 2.1], which may reduce the number of iterations and computation time (see Figure 4). Moreover, when $f$ and $\phi$ are $\mathcal{C}^{2}$, the pair $(f,\phi)$ is $L$-smad if and only if $L\nabla^{2}\phi(x)\preceq\nabla^{2}f(x)\preceq L\nabla^{2}\phi(x)$ holds for any $x\in C$. In order to achieve superior performance, it is recommended to choose a smaller $L$ and a $\phi$ that shares a similar structure with $f$. See, for more examples of the $L$-smad property, [Bolte-2018, Lemma 5.1][Bauschke2017-hg, Lemmas 7 and 8], [Takahashi, Proposition 24], [Mukkamala2019-mk, Proposition 2.1], [Takahashi2026-rv, Theorem 4.1], [Takahashi2023-uh, Theorem 1], and [Dragomir2021-rv, Propositions 2.1 and 2.3].
Input: $x^{0}\in\mathbb{R}^{n}$, $0<c_{1}<c_{2}<1$, $\lambda>0$
1
2for $k=0,1,2,\dots$ do
3   $y^{k}\leftarrow\tilde{\mathcal{T}}_{\lambda}(x^{k})$
4   $d^{k}\leftarrow y^{k}-x^{k}$
5   Compute $t_{k}$ such as $A_{k}(t_{k})<0$ and $W_{k}(t_{k})>0$ hold.
6
7  $x^{k+1}\leftarrow\left\{\begin{array}[]{ll}y^{k}&\text{if}\quad\Psi(y^{k})<\Psi(x^{k}+t_{k}d^{k}),\\ x^{k}+t_{k}d^{k}&\text{otherwise}.\end{array}\right$.
8
Algorithm 1 Approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search (ABPG-VMAW)
In the next section, we demonstrate that the search direction and step-size in the line search are well-defined and that the sequence of points generated by ABPG-VMAW globally converges to a stationary point.
4 Convergence Analysis
Throughout this section, we make the following assumption.
Assumption 4.1.
For a positive number $\sigma>0$, $\phi$ is $\sigma$-strongly convex on $C$.
Under Assumption 4.1, since $\tilde{\mathcal{T}}_{\lambda}(x)$ is strongly convex and closed, it has a unique minimizer. If $\phi$ is convex but not strongly convex, one can enforce strong convexity by adding the quadratic term $\frac{1}{2}\|\cdot\|^{2}$. When $g\equiv 0$, this modification does not affect the difficulty of the subproblems. When $g\not\equiv 0$, we are merely adding a quadratic and separable term $\frac{1}{2}\|\cdot\|^{2}$ to the subproblem that already contains a quadratic term. Therefore, the subproblems remain relatively easy to solve.
4.1 Properties of Proposed Algorithm
We first show the search direction property. More precisely, we prove that $d$ is a descent direction. The following inequality is a modified version of [Takahashi, Proposition 15].
Proposition 4.2 (Search direction property).
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. For any $x\in\operatorname{int}\operatorname{dom}\phi$, let $\xi\in\partial g(x)$. For any $\lambda>0$ and d=yxd=y-x defined by
$y=\tilde{\mathcal{T}}_{\lambda}(x)$ (4.1)
we have
$\langle\nabla f(x)+\xi,d\rangle\leq\langle\nabla f(x),d\rangle+g(x+d)-g(x)\leq-\frac{1}{\lambda}\langle\nabla^{2}\phi(x)d,d\rangle<0.$ (4.2)
Proof.
Since $g$ is convex, we have
$\langle\xi,y-x\rangle\leq g(y)-g(x),$ (4.3)
which implies
$\langle\nabla f(x)+\xi,d\rangle\leq\langle\nabla f(x),d\rangle+g(x+d)-g(x).$ (4.4)
From the first-order optimality condition of (4.1), we have
$-\nabla f(x)-\frac{1}{\lambda}\nabla^{2}\phi(x)(y-x)\in\partial(g+\delta_{\operatorname{cl}{C}})(y).$ (4.5)
Since $g$ is convex and $\delta_{\operatorname{cl}{C}}(x)=\delta_{\operatorname{cl}{C}}(y)=0$ from Lemma 3.3, it holds that
$g(y)-g(x)\leq-\left\langle\nabla f(x)+\frac{1}{\lambda}\nabla^{2}\phi(x)(y-x),y-x\right\rangle.$ (4.6)
Therefore, substituting $y\leftarrow x+d$ into the above inequality, we obtain
$\langle\nabla f(x),d\rangle+g(x+d)-g(x)$ $\leq\langle\nabla f(x),d\rangle-\left\langle\nabla f(x)+\frac{1}{\lambda}\nabla^{2}\phi(x)d,d\right\rangle$ (4.7)
$=-\frac{1}{\lambda}\langle\nabla^{2}\phi(x)d,d\rangle<0,$ (4.8)
where the last inequality holds because $\phi$ is strongly convex. ∎
When $t$ satisfies (3.6), we guarantee that the objective function value decreases.
Lemma 4.3 (Sufficient decrease property).
Suppose that Assumptions 3.1, 3.2, and 4.1 hold and that $t>0$ satisfies (3.6). For any $\lambda>0$, $x\in\operatorname{int}\operatorname{dom}\phi$ and d=yxd=y-x defined by (4.1), the following inequality holds:
$\Psi(x^{+})-\Psi(x)\leq-\frac{c_{1}t}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\leq 0,$ (4.9)
where
$x^{+}=\left\{\begin{array}[]{ll}y,&\text{if}\quad\Psi(y)<\Psi(x+td),\\ x+td,&\text{otherwise}.\end{array}\right$. (4.12)
Proof.
Let $\xi\in\partial g(x)$. Because (3.6) holds, $\delta_{\operatorname{cl}{C}}(x+td)=0$. From $y=x+d$, we have
$\Psi(x^{+})-\Psi(x)$ $\leq\Psi(x+td)-\Psi(x)$ (4.13)
$<c_{1}t\left(\langle\nabla f(x),d\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)$ (4.14)
$\leq-\frac{c_{1}t}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\leq 0,$ (4.15)
where the first inequality holds from (4.12), the second inequality holds from (3.6), and the last inequality holds from (4.2). ∎
The above lemma indicates that the objective function value is reduced at every step.
4.2 Global Subsequential Convergence
In this subsection, we discuss global subsequence convergence. In other words, we show that any accumulation point of a sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by ABPG-VMAW is a stationary point of (1.1). We use the limiting subdifferential and define the stationary point, inspired by Fermat’s rule [rockafellar2009variational, Theorem 10.1].
Definition 4.4.
A point $x^{*}\in\mathbb{R}^{n}$ is called a stationary point of $\Psi$ if
$0\in\nabla f(x^{*})+\partial(g+\delta_{\operatorname{cl}{C}})(x^{*}).$ (4.16)
Note that $\partial\delta_{\operatorname{cl}{C}}(x)=\{0\}$ if $x\in C$ because $C$ is open. When $x^{*}\in C$, $\nabla f(x^{*})+\partial(g+\delta_{\operatorname{cl}{C}})(x^{*})=\nabla f(x^{*})+\partial g(x^{*})$ from Assumption 3.1(iii). We make the following assumption.
Assumption 4.5.
  1. (i)
    The objective function $\Psi$ is level-bounded, i.e., for any $r\in\mathbb{R}$, lower level sets $\{x\in\mathbb{R}^{n}\,|\,\Psi(x)\leq r\}$ is bounded.
  2. (ii)
    The step-size $t_{k}>0$ at every $k$th iteration satisfies $A_{k}(t_{k})<0$ and $W_{k}(t_{k})>0$.
  3. (iii)
    The step-size $t_{k}>0$ is upper bounded, i.e., there exists $\bar{t}<\infty$ such that $t_{k}<\bar{t}$ holds for any $k\in\mathbb{N}$.
Assumption 4.5(i) is often assumed in nonsmooth optimization when the problem includes nonsmooth lower semicontinuous functions [Bolte-2018, Takahashi]. In fact, a lower semicontinuous, level-bounded, and proper function has a minimum [rockafellar2009variational, Theorem 1.9]. Assumption 4.5(ii) would often hold when the influence of $f$ is dominant compared to that of $g$. We will discuss this issue in more detail in Section A. Moreover, under Assumption 4.5(ii), Assumption 4.5(iii) always holds because $\Psi$ is bounded below from Assumption 3.1(iv) and the right-hand side of (3.6) is unbounded below. In this case, we have $\|t_{k}d^{k}\|\to 0$.
Lemma 4.6.
Suppose Assumptions 3.1, 3.2, 4.1, and 4.5 hold. Let $\{t_{k}\}_{k\in\mathbb{N}}$ and $\{x_{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW, $\bar{t}$ be a upper bound of the sequence $\{t_{k}\}_{k\in\mathbb{N}}$, and $\{d_{k}\}_{k\in\mathbb{N}}$ be a sequence of search directions in each iteration of ABPG-VMAW. It holds that
$\lim_{k\to\infty}\|t_{k}d^{k}\|=0.$ (4.17)
Proof.
Substituting $\leftarrow x^{k}, x+x$$^{+}\leftarrow x^{k+1}$, $d\leftarrow d^{k}$, and $t\leftarrow t_{k}$ into (4.9) in Lemma 4.3, we have
$0\leq\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x^{k})t_{k}d^{k},t_{k}d^{k}\rangle\leq\frac{c_{1}t_{k}\bar{t}}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\leq\bar{t}\left(\Psi(x^{k})-\Psi(x^{{k}+1})\right),$ (4.18)
where the second inequality holds because of $t_{k}^{2}\leq t_{k}\bar{t}$. Since $\phi$ is $\sigma$-strongly convex, the above inequality provides
$\frac{c_{1}\sigma}{2\lambda}\|t_{k}d^{k}\|^{2}\leq\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x^{k})t_{k}d^{k},t_{k}d^{k}\rangle\leq\bar{t}\left(\Psi(x^{k})-\Psi(x^{{k}+1})\right).$ (4.19)
Summing (4.19) from $k=0$ to $\infty$, we obtain
$\frac{c_{1}\sigma}{2\lambda}\sum_{k=0}^{\infty}\|t_{k}d^{k}\|^{2}\leq\bar{t}\sum_{k=0}^{\infty}\left(\Psi(x^{k})-\Psi(x^{{k}+1})\right).$ (4.20)
Using $\Psi^{*}\coloneqq\inf\Psi(x)>$-$\infty$ from Assumption 3.1(iv), we have
$\frac{c_{1}\sigma}{2\lambda}\sum_{k=0}^{\infty}\|t_{k}d^{k}\|^{2}$ $\leq\bar{t}\sum_{k=0}^{\infty}\left(\Psi(x^{k})-\Psi(x^{{k}+1})\right)$ (4.21)
$\leq\bar{t}(\Psi(x^{0})-\liminf_{N\to\infty}\Psi(x^{N}))$ (4.22)
$\leq\bar{t}(\Psi(x^{0})-\Psi^{*})<\infty,$ (4.23)
which implies $\lim_{{k}\to\infty}\|t_{k}d^{k}\|=0$. ∎
We establish the global subsequential convergence of ABPG-VMAW.
Theorem 4.7 (Global subsequential convergence).
Suppose that Assumptions 3.1, 3.2, 4.1, and 4.5 hold. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW. Then, the following statements hold:
  1. (i)
    The sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded.
  2. (ii)
    Any accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ is a stationary point of (1.1).
Proof.
(i) Since $\Psi(x^{k+1})\leq\Psi(x^{k})$ from Lemma 4.3 and $\Psi$ is level-bounded, the sequence of points $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded.
(ii) Substituting $x\leftarrow x^{k}$ and $y\leftarrow y^{k}$ into (4.5) with $y^{k}=x^{k}+d^{k}$ yields
$-\nabla f(x^{k})-\frac{1}{\lambda}\nabla^{2}\phi(x^{k})d^{k}\in\partial(g+\delta_{\operatorname{cl}{C}})(x^{k}+d^{k}).$ (4.24)
Since $g$ is convex, it follows that for $\xi^{k}\in\partial g(x^{k})=\partial(g+\delta_{\operatorname{cl}{C}})(x^{k})$ and $\nabla f(x^{k})-\frac{1}{\lambda}\nabla^{2}\phi(x^{k})d^{k}\in\partial(g+\delta_{\operatorname{cl}{C}})(x^{k}+d^{k})$,
$\left\langle-\nabla f(x^{k})-\frac{1}{\lambda}\nabla^{2}\phi(x^{k})d^{k}-\xi^{k},d^{k}\right\rangle\geq 0.$ (4.25)
This implies
$\langle-\nabla f(x^{k})-\xi^{k},d^{k}\rangle\geq\frac{1}{\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\geq\frac{\sigma}{\lambda}\|d^{k}\|^{2},$ (4.26)
where the last inequality holds because $\phi$ is $\sigma$-strongly convex. Let $\bar{x}\in\mathbb{R}^{n}$ be an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$ and let $\{x^{k_{j}}\}_{j\in\mathbb{N}}$ be a subsequence such that $x^{k_{j}}\to\bar{x}$ by the Bolzano–Weierstrass theorem. From (4.26) and the Cauchy–Schwarz inequality, we have
$\frac{\sigma}{\lambda}\|d^{k}\|^{2}\leq\langle-\nabla f(x^{k})-\xi^{k},d^{k}\rangle\leq\|\nabla f(x^{k})+\xi^{k}\|\|d^{k}\|.$ (4.27)
If $\|\nabla f(x^{k})+\xi^{k}\|=0$, then $x^{k}$ becomes a stationary point. We assume $\|\nabla f(x^{k})+\xi^{k}\|>0$. By the triangle inequality, we have $\|\nabla f(x^{k})+\xi^{k}\|\leq\|\nabla f(x^{k})\|+\|\xi^{k}\|$. Since the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, $\|\nabla f(x^{k})\|$ is bounded by the extreme value theorem and $\|\xi^{k}\|$ is also bounded (see, e.g.[Takahashi2022-ml, Theorem 1(ii)]). Thus, $\|\nabla f(x^{k})+\xi^{k}\|$ is bounded, i.e., due to (4.27), $d^{k}$ is also bounded. Thus, there exists a subsequence $\{d^{k_{j}}\}_{j\in\mathbb{N}}$ such that $d^{k_{j}}\to\bar{d}$ as $j\to\infty$ by the Bolzano–Weierstrass theorem. Then, by Lemma 4.6, the sequence $\{x^{k_{j}}+t_{k_{j}}d^{k_{j}}\}_{j\in\mathbb{N}}$ also converges to $\bar{x}$.
If $\liminf_{j\to\infty}t_{k_{j}}>0$, then it follows by (4.17) that $\lim_{j\to\infty}\|d^{k_{j}}\|=0$. Therefore, we only need to consider the case where $\liminf_{j\to\infty}t_{k_{j}}=\lim_{j\to\infty}t_{k_{j}}=0$. Let $\{\xi^{k_{j}}\}_{j\in\mathbb{N}}$ be a subsequence of $\xi^{k_{j}}\in\partial g(x^{k_{j}})$ so that $\xi^{k_{j}}\to\bar{\xi}$ as $j\to\infty$. Relabeling the indices again if necessary, we can choose the index set $\{k_{j}\}_{j\in\mathbb{N}}$ such that the sequences $\{x^{k_{j}}\}_{j\in\mathbb{N}}$, $\{d^{k_{j}}\}_{j\in\mathbb{N}}$, and $\{\xi^{k_{j}}\}_{j\in\mathbb{N}}$ converge to $\bar{x}$, $\bar{d}$, and $\bar{\xi}$, respectively.
From the condition (3.7), we have
$\langle\nabla f(x^{k_{j}}+t_{k_{j}}d^{k_{j}})+\xi^{k_{j}},d^{k_{j}}\rangle>c_{2}\langle\nabla f(x^{k_{j}})+\xi^{k_{j}},d^{k_{j}}\rangle.$ (4.28)
As $j\to\infty$, we have
$\langle\nabla f(\bar{x})+\bar{\xi},\bar{d}\rangle\geq c_{2}\langle\nabla f(\bar{x})+\bar{\xi},\bar{d}\rangle,$ (4.29)
which implies, due to $0<c_{2}<1$,
$\langle\nabla f(\bar{x})+\bar{\xi},\bar{d}\rangle\geq 0.$ (4.30)
Moreover, from (4.26), it holds that
$\langle\nabla f(x^{k_{j}})+\xi^{k_{j}},d^{k_{j}}\rangle\leq-\frac{\sigma}{\lambda}\|d^{k_{j}}\|^{2}\leq 0.$ (4.31)
Taking the limit as $j\to\infty$ and using (4.30), we have
$0\leq\langle\nabla f(\bar{x})+\bar{\xi},\bar{d}\rangle\leq-\frac{\sigma}{\lambda}\|\bar{d}\|^{2}\leq 0,$ (4.32)
which implies $d^{k_{j}}\to 0$. Because $f$ and $g$ are lower semicontinuous, from (4.24), we have
$0\in\nabla f(\bar{x})+\partial(g+\delta_{\operatorname{cl}{C}})(\bar{x}).$ (4.33)
We conclude that $\bar{x}$ is a stationary point. ∎
Assumption 4.8.
$\nabla f$ is Lipschitz continuous on any compact subset of $\mathbb{R}^{n}$.
Assumption 4.8 is weaker than the global Lipschitz continuity for $\nabla f$. Since $\phi$ is $\mathcal{C}^{2}$, $\nabla\phi$ is Lipschitz continuous on any compact subset of $\mathbb{R}^{n}$.
Lemma 4.9 (Lower bound of $t_{k}$).
Suppose Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Let $\{t_{k}\}_{k\in\mathbb{N}}$ be a sequence of points generated by ABPG-VMAW. For any $k\in\mathbb{N}$, there exists $\underline{t}>0$ such that $t_{k}>\underline{t}$ holds.
Proof.
From the condition (3.7), for $\xi^{k}\in\partial g(x^{k})$, we have
$\langle\nabla f(x^{k}+t_{k}d^{k})+\xi^{k},d^{k}\rangle>c_{2}\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle.$ (4.34)
There exists an $M_{1}>0$ such that the following inequality holds:
$M_{1}t_{k}\|d^{k}\|^{2}$ $\geq\langle\nabla f(x^{k}+t_{k}d^{k})-\nabla f(x^{k}),d^{k}\rangle$ (4.35)
$>c_{2}\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle-\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle$ (4.36)
$=-(1-c_{2})\langle\nabla f(x^{k})+\xi^{k},d^{k}\rangle,$ (4.37)
where the first inequality holds due to the Cauchy–Schwarz inequality and $\nabla f$ being Lipschitz continuous on any compact subset from Assumption 4.8, and the second inequality holds due to (4.34). Moreover, using the inequality (4.26), we have
$t_{k}\geq\frac{(1-c_{2})\langle-\nabla f(x^{k})-\xi^{k},d^{k}\rangle}{M_{1}\|d^{k}\|^{2}}\geq\frac{(1-c_{2})\sigma}{M_{1}\lambda}>0$ (4.38)
and therefore $\lim_{k\to\infty}t_{k}=\frac{(1-c_{2})\sigma}{M_{1}\lambda}>0$ holds. ∎
Proposition 4.10.
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW and $\underline{t}$ be a lower bound of $\{t_{k}\}_{k\in\mathbb{N}}$. Then, $\lim_{k\to\infty}\|d^{k}\|=0$ holds.
Proof.
Lemma 4.9 shows there exists a lower bound $\underline{t}:=\inf_{k}t_{k}>0$. From Lemma 4.3 and $t_{k}\in(\underline{t},\bar{t}]$, we have
$\Psi(x^{k+1})-\Psi(x^{k})$ $\leq-\frac{c_{1}t_{k}}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle$ (4.39)
$\leq-\frac{c_{1}\underline{t}}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle.$ (4.40)
Using the above inequality with the $\sigma$-strong convexity of $\phi$, we obtain
$\frac{c_{1}\sigma\underline{t}}{2\lambda}\|d^{k}\|^{2}\leq\frac{c_{1}\underline{t}}{2\lambda}\langle\nabla^{2}\phi(x^{k})d^{k},d^{k}\rangle\leq\Psi(x^{k})-\Psi(x^{k+1}).$ (4.41)
Summing this inequality from $k=1$ to $\infty$ and Assumption 3.1(iv), we have
$\frac{c_{1}\sigma\underline{t}}{2\lambda}\sum_{k=1}^{\infty}\|d^{k}\|^{2}\leq\Psi(x^{0})-\Psi^{*}<\infty,$ (4.42)
which implies $\lim_{k\to\infty}\|d^{k}\|=0$. ∎
Now, by using Proposition 4.10 and an argument similar to that of Lemma 4.6, we have $\|x^{k+1}-x^{k}\|\to 0$.
4.3 Global Convergence
Now, we show the global convergence of ABPG-VMAW. Before discussing global convergence, we have the following lemma.
Lemma 4.11.
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW. Then, the following statements hold:
  1. (i)
    There exist $\rho>0$ and $w^{k}\in\nabla f(y^{k})+\partial(g+\delta_{\operatorname{cl}{C}})(y^{k})$ such that
    $\|w^{k}\|\leq\rho\|x^{k+1}-x^{k}\|.$ (4.43)
  2. (ii)
    $\Psi\equiv\zeta$ on $\Omega$, where $\Omega$ is the set of accumulation points of $\{x^{k}\}_{k\in\mathbb{N}}$. Moreover, $\lim_{k\to\infty}\Psi(y^{k})=\Psi(\bar{x})$ for any $\bar{x}\in\Omega$.
Proof.
(i) Because we can define $\underline{t}=\min\left\{1,\frac{(1-c_{2})\sigma}{M_{1}\lambda}\right\}$ if necessary, without loss of generality, we assume $\underline{t}\in(0,1]$. Let $w^{k}\coloneqq\nabla f(y^{k})-\nabla f(x^{k})-\frac{1}{\lambda}\nabla^{2}\phi(x^{k})(y^{k}-x^{k})$. Using (4.24), we have $w^{k}\in\nabla f(y^{k})+\partial(g+\delta_{\operatorname{cl}{C}})(y^{k})$. There exists $M_{1}$ and $M_{2}>0$ such that, for $w^{k}$ and any $k\in\mathbb{N}$, it holds that
$\|w^{k}\|$ $\leq\|\nabla f(y^{k})-\nabla f(x^{k})\|+\frac{1}{\lambda}\|\nabla^{2}\phi(x^{k})(y^{k}-x^{k})\|$ (4.44)
$\leq M_{1}\|y^{k}-x^{k}\|+\frac{M_{2}}{\lambda}\|y^{k}-x^{k}\|$ (4.45)
$\leq\frac{M_{1}+M_{2}/\lambda}{\underline{t}}\|x^{k+1}-x^{k}\|,$ (4.46)
where the second inequality holds because of the Lipschitz continuity of $\nabla f$ and $\nabla\phi$ on compact subsets from Assumption 4.8 and Assumption 3.1(i), and the last inequality holds from line 1 in Algorithm 1.
(ii) Take any $\bar{x}\in\Omega$, i.e., $\{x^{k_{j}}\}_{j\in\mathbb{N}}$ such that $\lim_{j\to\infty}x^{k_{j}}=\bar{x}$. From Proposition 4.10, we can take $\{y^{k_{j}}\}_{j\in\mathbb{N}}$ such that $\lim_{j\to\infty}y^{k_{j}}=\bar{x}$ due to $d^{k_{j}}=y^{k_{j}}-x^{k_{j-1}}$. It follows from the definition of $y^{k}$ that
$\langle\nabla f(x^{k-1}),y^{k}-x^{k-1}\rangle$ $+g(y^{k})+\frac{1}{\lambda}\tilde{D}_{\phi}(y^{k},x^{k-1})$ (4.47)
$\leq\langle\nabla f(x^{k-1}),\bar{x}-x^{k-1}\rangle+g(\bar{x})+\frac{1}{\lambda}\tilde{D}_{\phi}(\bar{x},x^{k-1}),$ (4.48)
which is equivalent to
$g(y^{k})\leq\langle\nabla f(x^{k-1}),\bar{x}-y^{k}\rangle+g(\bar{x})+\frac{1}{\lambda}\tilde{D}_{\phi}(\bar{x},x^{k-1})-\frac{1}{\lambda}\tilde{D}_{\phi}(y^{k},x^{k-1}).$ (4.49)
Substituting $k$ for $k_{j}$ and letting $k\to\infty$, we obtain
$\limsup_{j\to\infty}g(x^{k_{j}})\leq g(\bar{x}).$ (4.50)
Using the continuity of $f$, we have $\limsup_{j\to\infty}\Psi(x^{k_{j}})\leq\Psi(\bar{x})$. In addition, $\Psi$ is lower semicontinuous from Assumption 3.1, $\Psi(\bar{x})\leq\liminf_{j\to\infty}\Psi(x^{k_{j}})$. Therefore, since $\bar{x}\in\Omega$ is arbitrary, $\lim_{j\to\infty}\Psi(x^{k_{j}})=\Psi(\bar{x})\equiv\zeta$. From line 1 on Algorithm 1, $\Psi(x^{k_{j}})\leq\Psi(y^{k_{j}})\leq\Psi(x^{k_{j-1}})$ implies $\lim_{j\to\infty}\Psi(y^{k_{j}})=\Psi(\bar{x})\equiv\zeta$. ∎
We establish that a sequence generated by ABPG-VMAW converges to a stationary point of (1.1).
Theorem 4.12 (Global convergence).
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Furthermore, suppose that $\Psi$ is a KL function. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW. Then, the following statements hold:
  1. (i)
    If $x^{k_{0}+k},y^{k_{0}+k-1}\in B(\bar{x},\rho)$ for some $k_{0}\in\mathbb{N}$, it holds that
    $2\|x^{k_{0}+k+1}-x^{k_{0}+k}\|\leq\|x^{k_{0}+k}-x^{k_{0}+k-1}\|+\chi_{k_{0}+k},$ (4.51)
    where $\chi_{k}=\frac{\rho_{2}}{\rho_{1}}[\psi(\Psi(x^{k})-\Psi(\bar{x}))-\psi(\Psi(x^{k+1})-\Psi(\bar{x}))]$.
  2. (ii)
    There exists $\bar{k}_{0}\in\mathbb{N}$ such that, for any $k\geq 1$, the following conditions hold:
    $x^{\bar{k}_{0}+k},y^{\bar{k}_{0}+k-1}\in B(\bar{x},\rho),$ (4.52)
    $\sum_{i=\bar{k}_{0}}^{\bar{k}_{0}+k}\|x^{i+1}-x^{i}\|+\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|\leq\|x^{\bar{k}_{0}+1}-x^{\bar{k}_{0}}\|+\chi_{\bar{k}_{0}+k}.$ (4.53)
  3. (iii)
    The sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to a stationary point of (1.1); moreover, $\sum_{k=0}^{\infty}\|x^{k+1}-x^{k}\|<\infty$.
Proof.
(i) Since $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded and $\Omega$ is the set of accumulation points of $\{x^{k}\}_{k\in\mathbb{N}}$ from Lemma 4.11(ii), we have $\lim_{k\to\infty}\operatorname{dist}(x^{k},\Omega)=0$, i.e.,
$\lim_{k\to\infty}\Psi(x^{k})=\Psi(\bar{x}).$ (4.54)
From Theorem 4.7, $\Omega$ is a subset of stationary points. Thus, if there exists an integer $\bar{k}\geq 0$ such that $\Psi(x^{k})=\Psi(\bar{x})$ holds for any $k\geq\bar{k}$, Lemma 4.3 implies $x^{\bar{k}+1}=x^{\bar{k}}$. A trivial induction shows that $\{x^{k}\}_{k\in\mathbb{N}}$ converges to a stationary point. Since $\{\Psi(x^{k})\}_{k\in\mathbb{N}}$ is a non-increasing sequence, (4.54) provides $\Psi(\bar{x})<\Psi(x^{k})$ for all $k\geq 0$. Again from (4.54), for any $v\in(0,+\infty]$, there exists an integer $k_{1}\geq 0$ such that, for all $k\geq k_{1}$, $\Psi(\bar{x})<\Psi(x^{k})<\Psi(\bar{x})+v$. From Lemma 4.11(ii), there exists an integer $k_{2}\geq 0$ such that, for all $k\geq k_{2}$, $\Psi(\bar{x})<\Psi(y^{k-1})<\Psi(\bar{x})+v$. Using this, the non-increase of $\{\Psi(x^{k})\}_{k\in\mathbb{N}}$, and line 1 in Algorithm 1, we have the following inequality for $k_{0}\geq\max\{k_{1},k_{2}\}$:
$\Psi(\bar{x})\leq\Psi(x^{k_{0}+k+1})<\Psi(x^{k_{0}+k})<\Psi(y^{k_{0}+k-1})<\Psi(\bar{x})+v,$ (4.55)
which implies $x^{k_{0}+k},y^{k_{0}+k-1}\in B(\bar{x},\rho)\cap\{z\in\mathbb{R}^{n}\,|\,\Psi(\bar{x})<\Psi(z)<\Psi(\bar{x})+v\}$. Here, by using Lemma 2.6 at $y^{k_{0}+k-1}$ and Lemma 4.11(i), we obtain
$\frac{1}{\rho_{2}\|x^{k_{0}+k}-x^{k_{0}+k-1}\|}\leq\frac{1}{\|w^{k_{0}+k-1}\|}\leq\psi^{\prime}(\Psi(y^{k_{0}+k-1})-\Psi(\bar{x}))\leq\psi^{\prime}(\Psi(x^{k_{0}+k})-\Psi(\bar{x})),$ (4.56)
where the last inequality holds from non-increase of $\psi^{\prime}$ due to concavity and $\Psi(y^{k_{0}+k-1})-\Psi(\bar{x})\geq\Psi(x^{k_{0}+k})-\Psi(\bar{x})$. Because $\psi$ is concave, it also holds that
$\psi(\Psi(x^{k_{0}+k})-\Psi(\bar{x}))-\psi(\Psi(x^{k_{0}+k+1})-\Psi(\bar{x}))$ (4.57)
$\quad\geq\psi^{\prime}(\Psi(x^{k_{0}+k})-\Psi(\bar{x}))(\Psi(x^{k_{0}+k})-\Psi(x^{k_{0}+k+1}))$ (4.58)
$\quad\geq\frac{\rho_{1}\|x^{k_{0}+k+1}-x^{k_{0}+k}\|^{2}}{\rho_{2}\|x^{k_{0}+k}-x^{k_{0}+k-1}\|},$ (4.59)
where the last inequality holds because of Lemma 4.3, $\sigma$-strongly convexity of $\phi$, and (4.56). By rearranging terms and letting $\chi_{k}=\frac{\rho_{2}}{\rho_{1}}[\psi(\Psi(x^{k})-\Psi(\bar{x}))-\psi(\Psi(x^{k+1})-\Psi(\bar{x}))]$, we obtain
$\|x^{k_{0}+k+1}-x^{k_{0}+k}\|^{2}\leq\chi_{k_{0}+k}\|x^{k_{0}+k}-x^{k_{0}+k-1}\|.$ (4.60)
Applying the arithmetic–geometric mean inequality yields
$2\|x^{k_{0}+k+1}-x^{k_{0}+k}\|\leq 2\sqrt{\chi_{k_{0}+k}\|x^{k_{0}+k}-x^{k_{0}+k-1}\|}\leq\chi_{k_{0}+k}+\|x^{k_{0}+k}-x^{k_{0}+k-1}\|.$ (4.61)
(ii) Without loss of generality, we assume that $\underline{t}\in(0,1]$ is the lower bound of $\{t_{k}\}_{k\in\mathbb{N}}$ (see also the proof of Lemma 4.11(i)). Let $\psi\in\Xi_{v}$. To establish (ii), we prove that there exists a sufficiently large integer $k_{0}$ such that
$\|\bar{x}-x^{k_{0}}\|+3\sqrt{\frac{\Psi(x^{k_{0}})-\Psi(\bar{x})}{\rho_{1}\underline{t}^{2}}}+\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{k_{0}})-\Psi(\bar{x}))<\rho,$ (4.62)
and then prove that $\|x^{k_{0}+k}-\bar{x}\|$ and $\|y^{k_{0}+k}-\bar{x}\|$ are bounded by the left-hand side of (4.62). Note that $k_{0}$ needs to be larger than $k_{1}$ and $k_{2}$ mentioned above.
From (4.54), there exists a nonnegative integer $k_{3}$ such that it holds for any $k\geq k_{3}$ that
$3\sqrt{\frac{\Psi(x^{k})-\Psi(\bar{x})}{\rho_{1}\underline{t}^{2}}}<\frac{\rho}{3}\quad\text{and}\quad\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{k})-\Psi(\bar{x}))$ $<\frac{\rho}{3}.$ (4.63)
Note that since $0<\underline{t}\leq 1$ for any $k\geq k_{3}$, it holds that
$3\sqrt{\frac{\Psi(x^{k})-\Psi(\bar{x})}{\rho_{1}}}<\frac{\rho}{3}.$ (4.64)
Since $\bar{x}$ is an accumulation point of the sequence $\{x^{k}\}_{k\in\mathbb{N}}$, there exists a nonnegative integer $k_{4}\geq 0$ such that $\|\bar{x}-x^{k}\|<\rho/3$ holds for any $k\geq k_{4}$. Using (4.63) and defining $\bar{k}_{0}\geq\max\{k_{1},k_{2},k_{3},k_{4}\}$, we have (4.62).
Using (4.62), we prove that (4.52) and (4.53) hold for any $k\geq 1$ by induction. For $k=1$, from (4.41) and $\Psi(x^{\bar{k}_{0}})-\Psi(x^{\bar{k}_{0}+1})<\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})$, it holds that
$\|x^{\bar{k}_{0}+1}-x^{\bar{k}_{0}}\|\leq\sqrt{\frac{\Psi(x^{\bar{k}_{0}})-\Psi(x^{\bar{k}_{0}+1})}{\rho_{1}}}\leq\sqrt{\frac{\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})}{\rho_{1}}}.$ (4.65)
Combining $\|\bar{x}-x^{\bar{k}_{0}}\|<\rho/3$ for $\bar{k}_{0}\geq\max\{k_{1},k_{2},k_{3},k_{4}\}$, (4.64), and (4.65), we have
$\|\bar{x}-x^{\bar{k}_{0}+1}\|\leq\|\bar{x}-x^{k_{0}}\|+\|x^{\bar{k}_{0}}-x^{\bar{k}_{0}+1}\|<\rho,$ (4.66)
which implies $x^{\bar{k}_{0}+1}\in B(\bar{x},\rho)$. Moreover, using a similar discussion and (4.63), we have
$\|\bar{x}-y^{\bar{k}_{0}}\|\leq\|\bar{x}-x^{\bar{k}_{0}}\|+\|x^{\bar{k}_{0}}-y^{\bar{k}_{0}}\|<\rho,$ (4.67)
i.e., $y^{\bar{k}_{0}}\in B(\bar{x},\rho)$. Due to $x^{\bar{k}_{0}+1},y^{\bar{k}_{0}}\in B(\bar{x},\rho)$ and (4.51), (4.52) and (4.53) hold for $k=1$.
Next, we suppose that (4.52) and (4.53) hold for $k\geq 1$. Since $\psi$ is positive and monotonically increasing, and $\{\Psi(x^{k})\}_{k\in\mathbb{N}}$ is non-increasing, we have
$\chi_{\bar{k}_{0}+k}\leq\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}+k})-\Psi(\bar{x}))\leq\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})).$ (4.68)
It holds that
$\|x^{\bar{k}_{0}+k+1}-\bar{x}\|$ $\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\sum_{i=\bar{k}_{0}}^{\bar{k}_{0}+k}\|x^{i+1}-x^{i}\|+\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|$ (4.69)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\|x^{\bar{k}_{0}}-x^{\bar{k}_{0}+1}\|+\chi_{\bar{k}_{0}+k}$ (4.70)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\sqrt{\frac{\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})}{\rho_{1}}}+\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x}))<\rho,$ (4.71)
where the first inequality holds from the triangle inequality and $\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|\geq 0$, the second inequality holds from the assumption (4.53), the third inequality holds from (4.41) and (4.68), and the last inequality holds from (4.62). Moreover, we have
$\quad\|y^{\bar{k}_{0}+k}-\bar{x}\|$ (4.72)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\|x^{\bar{k}_{0}}-x^{\bar{k}_{0}+1}\|+\sum_{i=\bar{k}_{0}}^{\bar{k}_{0}+k}\|x^{i+1}-x^{i}\|+\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|$ (4.73)
$\quad+\|y^{\bar{k}_{0}+k}-x^{\bar{k}_{0}+k}\|$ (4.74)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\|x^{\bar{k}_{0}}-x^{\bar{k}_{0}+1}\|+\chi_{\bar{k}_{0}+k}+\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|/\underline{t}$ (4.75)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+\sqrt{\frac{\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})}{\rho_{1}}}+\sqrt{\frac{\Psi(x^{\bar{k}_{0}+k})-\Psi(x^{\bar{k}_{0}+k+1})}{\rho_{1}\underline{t}^{2}}}+\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x}))$ (4.76)
$\leq\|x^{\bar{k}_{0}}-\bar{x}\|+2\sqrt{\frac{\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x})}{\rho_{1}\underline{t}^{2}}}+\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}})-\Psi(\bar{x}))<\rho,$ (4.77)
where the first inequality holds from the triangle inequality and $\|x^{\bar{k}_{0}+k+1}-x^{\bar{k}_{0}+k}\|\geq 0$, the second inequality holds from the assumption (4.53) and line 1 in Algorithm 1, the third inequality holds from (4.41) and (4.68), and the last inequality holds from (4.62). These imply $x^{\bar{k}_{0}+k+1}\in B(\bar{x},\rho)$ and $y^{\bar{k}_{0}+k}\in B(\bar{x},\rho)$, i.e., (4.52) holds. Using (4.51) and (4.53) for $k$, we have (4.53) for $k+1$. Therefore, (4.52) and (4.53) hold for all $k\geq 1$.
(iii) Finally, we establish global convergence. In this case, since
$\sum_{i=\bar{k}_{0}}^{\bar{k}_{0}+k}\|x^{i+1}-x^{i}\|\leq\|x^{\bar{k}_{0}+1}-x^{\bar{k}_{0}}\|+\frac{\rho_{2}}{\rho_{1}}\psi(\Psi(x^{\bar{k}_{0}+1})-\Psi(\bar{x}))$ (4.78)
holds for any $k\in\mathbb{N}$, we have $\sum_{i=\bar{k}_{0}}^{\infty}\|x^{i+1}-x^{i}\|<+\infty$, which implies that $\{x^{\bar{k}_{0}+k}\}_{k\in\mathbb{N}}$ converges to some $x^{*}$. Since $\bar{x}$ is an accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$, we have $x^{*}=\bar{x}$ from Theorem 4.7. ∎
Finally, we establish convergence rates, which are derived from $\sum_{k=0}^{\infty}\|x^{k+1}-x^{k}\|<+\infty$ in the same way as, e.g., [Bonettini_2017, Theorem 3], [Takahashi2022-ml, Theorem 4], and [Attouch2009-wf, Theorem 2].
Theorem 4.13 (Convergence rates).
Suppose that Assumptions 3.1, 3.2, 4.1, 4.5, and 4.8 hold. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by ABPG-VMAW and let $\bar{x}$ be a stationary point of (1.1). Suppose further that $\Psi$ is a KL function with $\psi$ in the KL inequality (2.5) taking the form $\psi(s)=cs^{1-\theta}$ for some $\theta\in[0,1)$ and $c>0$. Then, the following statements hold:
  1. (i)
    If $\theta=0$, then the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to $\bar{x}$ in a finite number of iterations;
  2. (ii)
    If $\theta\in(0,1/2]$, then there exist $c_{1}>0$ and $\eta\in[0,1)$ such that $\|x^{k}-\bar{x}\|<c_{1}\eta^{k}$;
  3. (iii)
    If $\theta\in(1/2,1)$, then there exists $c_{2}>0$ such that $\|x^{k}-\bar{x}\|<c_{2}k^{-\frac{1-\theta}{2\theta-1}}$.
5 Numerical Experiments
In this section, we conducted numerical experiments to examine the performance of our algorithm. All numerical experiments were performed in Python 3.9 on a MacBook Pro with an Apple M1 Max and 64GB LPDDR5 memory.
5.1 $\ell_{p}$-Regularized Least Squares Problem
We consider the sparse $\ell_{p}$-regularized least squares problem, where $p$ is slightly larger than 11[doi:10.1137/18M1194456, 7472557]:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|Ax-b\|^{2}+\frac{\theta_{p}}{p}\|x\|^{p}_{p},$ (5.1)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $\theta_{p}>0$. Let $g\equiv 0$. We also use $f$ and $\phi$ given by
$f(x)=\frac{1}{2}\|Ax-b\|^{2}+\frac{\theta_{p}}{p}\|x\|^{p}_{p},\quad\text{and}\quad\phi(x)=\frac{1}{2}\|x\|^{2}+\frac{1}{p}\|x\|^{p}_{p}.$ (5.2)
Note that $f$ and $\phi$ are $\mathcal{C}^{1}$ if $p>1$ while $\nabla f$ and $\nabla\phi$ are not globally Lipschitz continuous. Although we can choose any $\lambda>0$, we use $\lambda$ given by $\lambda<1/L$ if $(f,\phi)$ is $L$-smad (see, for more details, Remark 3.5). Note that our algorithm does not require the $L$-smad property.
Proposition 5.1 (The $L$-smad property of $(f,\phi)$ [Takahashi, Proposition 24]).
Let $f$ and $\phi$ be as defined above. Then, for any $L>0$ satisfying
$L\geq\lambda_{\max}(A^{\top}A)+\theta_{p},$ (5.3)
the functions $L\phi-f$ and $L\phi+f$ are convex on $\mathbb{R}^{n}$, i.e., the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$.
The subproblem of BPG cannot be solved in closed form if $p>1$ because its optimality condition is a (p1)(p-1)th polynomial equation. On the other hand, $\nabla^{2}\phi(x)=I+(p-1)\operatorname{diag}(|x|^{p-2})$ is a diagonal matrix and $\tilde{\mathcal{T}}_{\lambda}(x)$ can be solved in closed form even if $g\not\equiv 0$ [Takahashi, Remark 25].
(a) Objective function values ($p=1.2$)
(b) Accuracy ($p=1.2$)
Figure 1: Comparison with ABPG-VMAW (blue), ABPG (orange), PG (green), and PGL (red) on the $\ell_{p}$ regularized least squares problem (5.1)
We compare ABPG-VMAW with ABPG [Takahashi], the proximal gradient algorithm (PG) with a constant step-size, and PG with line search (PGL). We set $c_{1}=0.99$, $c_{2}=0.999$, $\mu=0.9$, and $\eta=2$ for ABPG-VMAW and $c_{1}=0.99$ and $\delta=0.9$ for ABPG. Although $\nabla f$ is not Lipschitz continuous, PG uses the step-size $1/L$ given by (5.3). Note that PG does not guarantee global convergence. PGL searches $\lambda_{k}>0$ satisfying the descent lemma and uses the initial step-size $\lambda_{0}=1/L$ given by (5.3[doi:10.1137/1.9781611974997, p.283]. The initial point $x^{0}\in\mathbb{R}^{n}$ is generated from an i.i.d. normal distribution. The maximum number of iterations is 1000. The terminal condition is $\|x^{k}-x^{k-1}\|\leq 10^{-8}$.
(a) Objective function values ($p=1.2$)
(b) Accuracy ($p=1.2$)
Figure 2: Comparison with ABPG-VMAW (blue), ABPG (orange), PG (green), and PGL (red) on the $\ell_{p}$ regularized least squares problem (5.1)
The problem setting is as follows. We generate the matrix $A\in\mathbb{R}^{n\times m}$ and the ground truth $x^{*}\in\mathbb{R}^{n}$, which has 10% nonzero elements, from i.i.d. normal distribution. We set $b=Ax^{*}$. For $(n,m)=(1000,700)$, $p=1.2$, and $\theta_{p}=0.1$, Figure 1 shows the objective function value $\Psi(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale and Figure 2 shows those on the time axis. When $p=1.2$, the gradient of $\|x\|^{p}_{p}$ is not Lipschitz continuous on $(-1,1)^{n}$. This is why PG and PGL are not guaranteed to converge to a stationary point in this setting. According to Figures 1, 1, 2, and 2, when $p=1.2$, only ABPG-VMAW and ABPG converge within 1000 iterations, while PG and PGL do not satisfy the stopping condition. In particular, ABPG-VMAW meets the stopping condition in fewer than 200 iterations, which is significantly fewer than ABPG, which requires over 800 iterations.
Next, we show the average performance of the four methods—ABPG-VMAW, ABPG, PG, and PGL—on the $\ell_{p}$-regularized least squares problem. Specifically, we selected combinations of $m$ and $n$ from the set $\{100,200\}\times\{1000,2000,5000\}$. For each combination, we generated 100 random instances: in each instance, we drew an $m\times n$ matrix $A$ and a ground truth vector $x^{*}\in\mathbb{R}^{n}$ with $10\%$ nonzero entries from an i.i.d. normal distribution. For each generated instance, we set $b=Ax^{*}$, $p=1.2$, and $\theta_{p}=0.1$. Table 1 presents the average performance, including the number of iterations, the accuracy of the recovered point, the objective values, and computation time, across 100 different instances. ABPG-VMAW outperformed ABPG, PG, and PGL. Moreover, ABPG-VMAW converged in fewer iterations and in a shorter amount of time than ABPG, PG, and PGL.
Figure 3 shows which of $y^{k}$ and $x^{k}+t_{k}d^{k}$ in line 1 of Algorithm 1 is selected in ABPG-VMAW. The red plot represents the values of $t_{k}$ that are actually adopted, showing that $x^{k}+t_{k}d^{k}$ is selected. Figure 3 indicates that the proposed method allows larger step-sizes, which may account for its convergence in fewer iterations than ABPG. Moreover, for ABPG-VMAW, we varied $\lambda$ by scaling $L$ by each value in $\{0.01,0.05,0.1,0.2,0.5,0.8,0.9,1.0,1.1,1.2,2.0,5.0,10.0,100\}$, where $L$ given by (5.3) and compared the results. The maximum number of iterations is 15000. The results are shown in Figure 4. When the value around $L$ is chosen, both the number of iterations and the computation time are reduced.
Figure 3: Trace of step-size parameter $t_{k}$ over iterations
Figure 4: Comparative analysis of algorithmic performance with varying $L$-smad parameter scaling
Table 1: Average number of iterations, objective function value, accuracy, and CPU time for ABPG-VMAW, ABPG, PG, and PGL using random instances of the $\ell_{p}$-regularized least squares problem (5.1) ($100$ instances for $m=100$, and $10$ instances for $m=1000$)
$m$ $n$ algorithm iteration obj acc time
100 1500 ABPG-VMAW 125 0.494 0.993 0.227
ABPG 847 0.494 0.993 1.022
PG 1000 6.218 1.068 0.624
PGL 1000 17.968 4.385 0.686
3000 ABPG-VMAW 219 0.503 0.997 1.623
ABPG 980 0.503 0.997 3.750
PG 1000 6.412 1.039 3.174
PGL 1000 43.325 7.374 3.285
5000 ABPG-VMAW 160 0.492 0.999 2.072
ABPG 1000 0.492 0.999 6.497
PG 1000 6.372 1.025 5.618
PGL 1000 79.301 10.345 5.799
1000 1500 ABPG-VMAW 108 0.496 1.000 0.397
ABPG 587 0.496 1.000 3.473
PG 1000 28.745 2.067 0.777
PGL 1000 9.874 2.326 1.037
3000 ABPG-VMAW 44 0.512 1.000 0.548
ABPG 504 0.512 1.000 7.251
PG 1000 36.533 1.958 3.775
PGL 1000 24.393 3.981 4.343
5000 ABPG-VMAW 47 0.498 1.000 1.025
ABPG 434 0.498 1.000 9.925
PG 1000 43.081 1.875 6.493
PGL 1000 45.076 5.704 7.414
5.2 Nonnegative Linear Inverse Problem
Given a nonnegative matrix $A\in\mathbb{R}^{m\times n}_{+}$ and a nonnegative vector $b\in\mathbb{R}^{m}_{+}$, the goal of nonnegative linear inverse problems is to recover a signal $x\in\mathbb{R}^{n}_{+}$ such that $Ax\simeq b$. Nonnegative linear inverse problems have been studied in image deblurring [Bertero2009-jq] and positron emission tomography [Vardi1985-de], as well as in optimization [Bauschke2017-hg, Takahashi]. To achieve the goal of nonnegative linear inverse problems, we focus on the convex optimization problem given by
$\min_{x\in\mathbb{R}^{n}_{+}}\quad D_{\rm{KL}}(Ax+b)+\theta_{1}\|x\|_{1},$ (5.4)
where the Kullback–Leibler divergence is defined as follows:
$D_{\rm{KL}}(x,y)=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right).$ (5.5)
Let $f(x)=D_{\rm{KL}}(Ax,b)$ and $g(x)=\theta_{1}\|x\|_{1}$. We use $\phi_{0}(x)=\sum_{i=1}^{n}x_{i}\log x_{i}$ as the kernel generating distance for BPG and $\phi_{1}(x)=\phi_{0}+\frac{1}{2}\|x\|^{2}$ as the kernel generating distance for our algorithm and ABPG. In this case, we also define $C=\operatorname{int}\operatorname{dom}\phi_{0}=\operatorname{int}\operatorname{dom}\phi_{1}=\mathbb{R}^{n}_{+}$. When $\sum_{i=1}^{m}a_{ij}=1$, the pair $(f,\phi_{0})$ is $1$-smad [Bauschke2017-hg] and the pair $(f,\phi_{1})$ is also $1$-smad [Takahashi]. We compare ABPG-VMAW with ABPG [Takahashi], PGL, and BPG. Those subproblems can be solved in closed form.
The problem setting is as follows. We generate the matrix $A\in\mathbb{R}^{m\times n}$ and the ground truth $x^{*}\in\mathbb{R}^{n}$, which has 5% nonzero elements, from an i.i.d. normal distribution. We set $b=Ax^{*}$. For $(n,m)=(200,500)$ and $\theta_{1}=0.05$, Figure 5 shows the objective function value $\Psi(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale and Figure 6 shows those on the time axis.
Under this condition, ABPG-VMAW outperforms the other three methods in terms of the reduction in the objective function value per iteration and per unit time. It is also observed that the objective function values obtained after 1000 iterations are comparable across all methods. Notably, the error with respect to the true value is significantly smaller for ABPG-VMAW than for the other three methods.
(a) Objective function values
(b) Accuracy
Figure 5: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the nonnegative linear inverse Problem (5.4)
(a) Objective function values
(b) Accuracy
Figure 6: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the nonnegative linear inverse problem (5.4)
5.3 Phase Retrieval
We consider phase retrieval, i.e., recovering a signal $x\in\mathbb{R}^{n}$ such that $|\langle a_{i},x\rangle|^{2}\simeq b_{i}$ for $i=1,\ldots,m$, where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b_{i}\in\mathbb{R}$ is a observed magnitude. Phase retrieval has been studied in many applications, such as image processing [Candes2015-mk] and X-ray crystallography [Patterson1934-sg, Patterson1944-nm] as well as optimization [Bolte-2018, Takahashi2022-ml, Takahashi2025-py]. We address the following nonconvex optimization problem:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}.$ (5.6)
Let $f(x)=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2}$. When we use $\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2}$, $(f,\phi)$ is $L$-smad for any $L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$ (see [Bolte-2018, Lemma 5.1]). We compare ABPG-VMAW with ABPG, PGL, and BPG. Those subproblems can be solved in closed form (see [Bolte-2018, Proposition 5.1] for the subproblem of BPG). We set $\lambda=1/L$, where $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)$.
The problem setting is as follows. We generate $a_{i}\in\mathbb{R}^{n}$, $i=1,\ldots,m$, and the ground truth $x^{*}\in\mathbb{R}^{n}$ from an i.i.d. normal distribution. We set $b_{i}=|\langle a_{i},x^{*}\rangle|^{2}$. For $(n,m)=(200,1000)$, Figure 7 shows the objective function value $f(x^{k})$ and the accuracy $\|x^{k}-x^{*}\|$ at each iteration on a logarithmic scale, and Figure 8 shows those on the time axis.
Under this condition, ABPG-VMAW outperforms the other three methods in terms of the reduction in the objective function value per iteration and per unit time. ABPG-VMAW allows large step-sizes while ABPG and BPG use $\lambda=1/L$, which can be small, and PGL would estimate small step-sizes without Lipschitz continuous gradients. Moreover, the error with the ground truth is significantly smaller for ABPG-VMAW than for the other three methods.
(a) Objective function values
(b) Accuracy
Figure 7: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the phase retrieval (5.6)
(a) Objective function values
(b) Accuracy
Figure 8: Comparison with ABPG-VMAW (blue), ABPG (orange), PGL (green), and BPG (purple) on the phase retrieval (5.6)
6 Conclusion
In this paper, we propose the approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search (ABPG-VMAW) for composite nonconvex optimization problems. Our line search condition allows a larger step-size than existing algorithms. We have established global subsequential convergence under standard assumptions. Moreover, under the KL property, we have proved global convergence to a stationary point even when $g\not\equiv 0$. To the best of our knowledge, this is the first global convergence result for ABPG-type algorithms in this setting. Moreover, our numerical experiments on $\ell_{p}$ regularized least squares problems, nonnegative linear inverse problems, and phase retrieval have shown that ABPG-VMAW outperforms ABPG and proximal gradient algorithms.
On the other hand, our line search procedure would not be well-defined when the objective function is dominated by $g$ rather than by $f$ (in practice, this case is rare when $g$ is a regularizer). Although we establish that our line search is well-defined when $g\equiv 0$ in Section A, it is important to prove this in the general case $g\not\equiv 0$.
Declarations
Funding: This project has been funded by the Japan Society for the Promotion of Science (JSPS) and the Nakajima foundation; JSPS KAKENHI Grant Number JP23K19953 and JP25K21156; JSPS KAKENHI Grant Number JP23H03351.
Conflict of interest: The authors have no competing interests to declare that are relevant to the content of this article.
Data availability: The datasets generated during and/or analyzed during the current study are available in the GitHub repository, https://github.com/ShotaTakahashi/ApproximateBPG.
Appendix A Appendix: Implementation of Line Search
In order to obtain a step-size $t_{k}$ satisfying both (3.6) and (3.7), we adopt a bisection method for the line search procedure in Algorithm 2.
Input: Functions $f$, $g$, $\phi$ and $\lambda\in\mathbb{R}$
Output: Step-size $t$
1 Procedure $\textnormal{{line\_search}}_{k}(f,g,\phi,\lambda)$
2   Choose $0<c_{1}<c_{2}<1$ and $0<\mu<1<\eta$
3   $q_{1}\leftarrow 1$
4
5  if $A_{k}(q_{1})\geq 0$ then
6     while $A_{k}(q_{1})\geq 0$ do
7       $q_{2}\leftarrow q_{1}$
8       $q_{1}\leftarrow\mu q_{1}$
9
10
11       else
12        while $A_{k}(q_{1})<0$ do
13           $q_{2}\leftarrow q_{1}$
14           $q_{1}\leftarrow\eta q_{1}$
15
16
17           $\alpha\leftarrow\min\{q_{1},q_{2}\}$
18           $\beta\leftarrow\max\{q_{1},q_{2}\}$
19           $t\leftarrow(\alpha+\beta)/2$
20
21          loop
22             if $A_{k}(t)\geq 0$ then
23               $\beta\leftarrow t$
24
25               else if $W_{k}(t)\leq 0$ then
26                 $\alpha\leftarrow t$
27
28                 else
29                   return $t$
30
31                   $t\leftarrow(\alpha+\beta)/2$
32
33
Algorithm 2 Variable Metric Armijo–Wolfe Line Search
A.1 Special Case: $g\equiv 0$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$
We consider the special case $g\equiv 0$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$, i.e., $\Psi\equiv f$. $\ell_{p}$ regularized least squares problems in Section 5.1 used this setting. We have Armijo–Wolfe conditions for $x\in\operatorname{int}\operatorname{dom}\phi$ and $d\in\mathbb{R}^{n}$ as follows:
$A(t)$ $=f(x+td)-f(x)-c_{1}t\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)<0,$ (A.1)
$W(t)$ $=\langle\nabla f(x+td),d\rangle-c_{2}\langle\nabla f(x),d\rangle>0.$ (A.2)
We prove that (A.1) and (A.2) are well-defined, i.e., there exists a number $t$ such that (A.1) and (A.2) hold simultaneously.
Lemma A.1.
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. Let $\lambda>0$, $0<c_{1}<c_{2}<1$, and $x\in\operatorname{int}\operatorname{dom}\phi$, and let d=yxd=y-x be defined by (4.1). There exists a pair of positive numbers $(t_{\beta},t_{\alpha})$, where $t_{\beta}>t_{\alpha}$, such that $A(t)<0$ holds for any $t\in[0,t_{\alpha})$ and $A(t)\geq 0$ holds for any $t>t_{\beta}$.
Proof.
Differentiating $A(t)$ with respect to $t$ , we obtain
$A^{\prime}(t)=\langle\nabla f(x+td),d\rangle-c_{1}\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)$ (A.3)
and substituting $t=0$ yields
$A^{\prime}(0)=(1-c_{1})\langle\nabla f(x),d\rangle-\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle<0,$ (A.4)
where the last inequality holds from Proposition 4.2 and $c_{1}<1$. Combining $A(0)=0$, it holds that there exists a positive number $t_{\alpha}$ such that $A(t)<0$ for any $t\in[0,t_{\alpha})$.
Next, we show the existence of $t_{\beta}$. Since $c(t)=f(x+td)$ is bounded below from Assumption 3.1(iv) and $f(x)+c_{1}t\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)\to$-$\infty$ as $t\to\infty$, there exists a positive number $t_{\beta}$ such that for any $t>t_{\beta}$ the following inequality holds:
$f(x+td)\geq f(x)+c_{1}t\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right),$ (A.5)
which implies $A(t)\geq 0$. Note that, from the definition of $t_{\alpha}$, we have $t_{\alpha}<t_{\beta}$. ∎
Lemma A.2.
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. Let a pair of positive numbers $(t_{\alpha},t_{\beta})$, where $t_{\beta}>t_{\alpha}$, such that $A(t_{\alpha})<0$ and $A(t_{\beta})\geq 0$ hold. There exists a nonempty interval $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$ in $t_{\alpha},t_{\beta}]$ such that (A.1) and (A.2) hold.
Proof.
Since $A(t_{\alpha})<0$ holds, we can define $t^{*}$ by
$t^{*}\coloneqq\sup\left\{t\in[t_{\alpha},t_{\beta}]\,\middle|\,\forall s\in[t_{\alpha},t],\langle\nabla f(x+sd),d\rangle\leq c_{2}\langle\nabla f(x),d\rangle\right\}.$ (A.6)
Then, $\langle\nabla f(x+td),d\rangle\leq c_{2}\langle\nabla f(x),d\rangle$ holds almost everywhere on the interval $t_{\alpha},t^{*}]$, and therefore we obtain
$f(x+t^{*}d)-f(x+t_{\alpha}d)$ $=\int_{t_{\alpha}}^{t^{*}}\langle\nabla f(x+td),d\rangle dt$ (A.7)
$\leq\int_{t_{\alpha}}^{t^{*}}c_{2}\langle\nabla f(x),d\rangle dt$ (A.8)
$=c_{2}(t^{*}-t_{\alpha})\langle\nabla f(x),d\rangle$ (A.9)
$<c_{1}(t^{*}-t_{\alpha})\langle\nabla f(x),d\rangle,$ (A.10)
where the first inequality holds from $\langle\nabla f(x+td),d\rangle\leq c_{2}\langle\nabla f(x),d\rangle$, and the last inequality holds from $c_{1}<c_{2}$ and Proposition 4.2. By adding $\frac{c_{1}t^{*}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\leq-\frac{c_{1}t_{\alpha}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle$ and rearranging the terms, we obtain
$\quad f(x+t^{*}d)-c_{1}t^{*}\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)$ (A.11)
$<f(x+t_{\alpha}d)-c_{1}t_{\alpha}\left(\langle\nabla f(x),d\rangle+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right),$ (A.12)
which implies
$A(t^{*})<A(t_{\alpha})<0.$ (A.13)
From the continuity of $A(t)$, there exists a positive number $\Delta$ such that $A(t)<0$ holds for any $t$ in $t^{*},t^{*}+\Delta]$, and from the definition of $t^{*}$ there exists a nonempty subset $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$ of $t^{*},t^{*}+\delta]$ such that $W(t)>0$ always holds i.e., (A.1) and (A.2) hold simultaneously on $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$. ∎
Lemma A.2 ensures that when the sign of $A(t)$ changes from negative to positive at two points, an Armijo–Wolfe step-size exists between those two points.
Next, using Lemmas A.1 and A.2, we show the well-definedness of the line search procedure i.e., that the line search procedure terminates in a finite number of steps. Its proof is almost the same as [Lewis2013, Theorem 4.7].
Theorem A.3 (Well-definedness of the line search procedure).
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. Whenever the second loop of the line search procedure in an iteration terminates, the final trial step $t$ is an Armijo–Wolfe step, provided that $\lambda$ is small enough. If, on the other hand, the line search procedure does not terminate, then it eventually generates a nested sequence of finite intervals $\alpha,\beta]$, halving in length at each iteration, and each containing a set of nonzero measure of Armijo–Wolfe steps. These intervals converge to a step $t_{0}>0$ such that
$A(t_{0})=0\quad\text{and}\quad W(t_{0})>0$ (A.14)
hold, i.e., $t_{0}$ is an Armijo–Wolfe step.
Remark A.4.
When $g\not\equiv 0$, Algorithm 2 would not terminate in finite steps to obtain $t_{k}$ such that (3.6) and (3.7) hold. For example, the influence of $g$ plays a principal role in determining the overall behavior of the objective function. However, this case is rare in practice because $g$ is often a regularizer. In fact, Algorithm 2 succeeds in obtaining $t_{k}$ satisfying (3.6) and (3.7) in Section 5.2.
Remark A.5.
Let $x\in\operatorname{int}\operatorname{dom}\phi$, $d\in\mathbb{R}^{n}$, and $t\in\mathbb{R}_{+}$. Suppose $d\neq 0$ because $\langle\nabla f(x+td)+\xi,d\rangle\geq c_{2}\langle\nabla f(x)+\xi,d\rangle$ always holds when $d=0$. We assume that $\phi$ is $\sigma$-strongly convex (see Assumption 4.1) and $g$ is $\kappa$-Lipschitz continuous with $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$, i.e., $\max_{\xi\in\partial g(x)}\|\xi\|\leq\kappa$ for any $x\in\operatorname{dom}g$. When $g$ is strictly continuous on $\operatorname{dom}g$ (e.g., locally Lipschitz continuous functions are strictly continuous), its subdifferential at any point of $\operatorname{dom}g$ is bounded [rockafellar2009variational, Theorem 9.13]. Using this and letting $\xi\in\partial g(x)$ and $\xi^{+}\in\partial g(x+td)$, we have
$0\leq\langle\xi^{+}-\xi,td\rangle\leq(\|\xi^{+}\|+\|\xi\|)\|td\|\leq 2\kappa t\|d\|,$ (A.15)
which implies
$\langle\xi^{+}-\xi,d\rangle\leq 2\kappa\|d\|\leq\frac{c_{1}\sigma}{2\lambda}\|d\|^{2}\leq\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle,$ (A.16)
where the second inequality holds from $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$, and the last inequality holds because of the strong convexity of $\phi$.
Suppose that $t$ satisfies the following Armijo condition:
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $\geq c_{1}\left(\langle\nabla f(x),d\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right).$ (A.17)
The existence of such $t$ is assured by a well-known discussion, e.g.[Nocedal2006-sb], and the mean-value theorem. We obtain
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $\geq c_{1}\left(\langle\nabla f(x),d\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)$ (A.18)
$\geq c_{1}\langle\nabla f(x)+\xi,d\rangle+\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle,$ (A.19)
where the last inequality holds from the convexity of $g$. On the other hand, using (A.16), we also have
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $=\langle\nabla f(x+td)+\xi,d\rangle+\langle\xi^{+}-\xi,d\rangle$ (A.20)
$\leq\langle\nabla f(x+td)+\xi,d\rangle+\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle.$ (A.21)
Combining these inequalities and using $\langle\nabla f(x)+\xi,d\rangle<0$ from (4.2) and $c_{1}<c_{2}$ implies
$\langle\nabla f(x+td)+\xi,d\rangle\geq c_{1}\langle\nabla f(x)+\xi,d\rangle\geq c_{2}\langle\nabla f(x)+\xi,d\rangle.$ (A.22)
Therefore, our line search procedure is well-defined. However, $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$ does not holds when $g$ is a dominance term, i.e., $\kappa>\frac{c_{1}\sigma}{4\lambda}\|d\|$ holds. In this case, our line search procedure might not be well-defined.